# NPHardEval4V: A Dynamic Reasoning Benchmark of Multimodal Large Language Models

### v2

@@ 3.1 NPHardEval Benchmark and Our Transformation
The figures in algorithmic questions such as SPP, GCP__\__D, TSP, and TSP__\__D are transformed from text similarly: these figures include graphs containing nodes and edges; in the case of TSP and TSP__\__D, the weights on the edges are also included.
Linear Data Transformation
Similarly, linear data problems like the Knapsack Problem (see Figure 3) are presented with both a prompt and a visual representation of items with associated weights and values. In order to provide a visual representation of the Knapsack problem, we use Python to create blocks with sizes corresponding to the weights of the items as described in the textual problem description. Specifically, each item is assigned a unique identifier (id) and a corresponding weight, which is used to determine the size of the block representing that item. The MLLMs are tasked with maximizing value within the constraints provided.
@@
For algorithmic problems such as EDP and SAS, we do not employ a specialized visualization method. Instead, we simply display the relevant string of characters or string of numbers in the image.
To summarize, we employ various visualization functions to transform textual descriptions into visual representations for different algorithmic problems. In the prompt, we provide a combination of textual and visual information to the model, including an instructional prompt and a general introduction of the algorithm, along with an image that represents the specific problem to be solved. This approach is demonstrated in the previous examples for GCP and KSP. By providing both textual and visual information, we aim to evaluate the reasoning abilities of MLLMs in handling complex problems that require both visual and textual understanding.
4 Experimental Setting
To address the outlined research questions, our experimental setting systematically assesses the performance of various MLLMs in recognition and reasoning tasks and investigates the impact of vision and text inputs on MLLMs’ performance. The models we test on are presented in Figure 2. We detail the experimental setup corresponding to each research question, describing the models evaluated, the nature of the tasks, the methodology of the performance assessment, and the structure of the benchmark.
@@ 4 Experimental Setting
GPT-4V GPT-4V Unknown Close source
Gemini 1.0 Pro GeminiPro Unknown Close source
CogVLM wang2023cogvlm 17 Open source
LLaVA-1.5-13B liu2024visual 13 Open source
BLIP-2 FLAN-T5-XXL li2023blip 11 Open source
Fuyu-8B fuyu-8b 8 Open source
Otter li2023mimic 8 Open source
Qwen-VL-7B bai2023qwen 7 Open source
Kosmos2 peng2023kosmos 1.5 Open source
Table 2: MLLMs metadata
4.1 Recognition and Reasoning Performance Evaluation
@@
In the initial experiment, the objective is to assess the performance of Multi-Modal Large Language Models (MLLMs) in terms of recognition and reasoning across a range of benchmark problems. To this end, zero-shot prompts are utilized to evaluate the inherent recognition and reasoning capabilities of each model. The performance of the models is then analyzed and compared across the benchmark problems to draw conclusions about their relative strengths and weaknesses in recognition and reasoning tasks.
Recognition Experiment
@@
In order to precisely assess the capacity for reasoning in Multi-Modal Large Language Models (MLLMs), it is essential to first comprehend their proficiency in image recognition, as accurate reasoning can only be established upon correct recognition. Therefore, an evaluation of recognition capabilities is initially conducted. For each question, both a visual representation of the description and a textual representation are provided. Subsequently, the MLLM is presented with both the visual and textual representations and is asked to determine whether they correspond to the same question. To mitigate the influence of randomness, the model is prompted five times, each time with a distinct random seed. The mean and variance of the model’s responses are then calculated, and it is concluded that the model demonstrates recognition capabilities if it correctly identifies the correspondence between the visual and textual representations more than half of the time, i.e., answering ’yes’ three or more times. As an preprocess for analyzing reasoning ability, recognition experiment reuslts and analysis are provided in Appendix B
@@
Reasoning Experiment 1, the Default Setup (Vision with Instructional Text)
@@
This experiment evaluates the overall performance of MLLM on a benchmark consisting of various tasks. In the setting, a textual prompt is provided that includes a general introduction to the question and the format in which the answer should be given. Subsequently, an image is presented that pertains to the specific question being asked. The models are then tasked with processing both the textual and visual information in order to generate accurate responses. The performance of the MLLMs is evaluated based on their ability to correctly answer the questions, taking into account both their recognition and reasoning capabilities. The results of this experiment provide insights into the overall effectiveness of the models in handling multimodal tasks and their potential applications in real-world scenarios.
@@
Figure 4: Image of an Short Path Problem example
In order to accurately evaluate the reasoning performance, it is necessary to isolate the specific aspect of reasoning ability from other factors that may influence the overall performance: we employ a filtering process that removes datapoints where the model fails to recognize the given input or fails to provide a parsable result, which indicates a failure in instruction-following. After filtering out these two factors, we can obtain a more accurate picture of the MLLMs’ pure reasoning ability.
4.2 Ablation on Multimodal prompts
The goal of this experiment is to investigate the impact of prompt modality on the reasoning performance of MLLMs. Specifically, we aim to assess whether the inclusion of visual or textual inputs can enhance the models’ problem-solving abilities. To achieve this, we conduct an ablation study, varying the modality of the prompts provided to the MLLMs by pure textual prompt and textual + visual prompt. By comparing the models’ performance across different prompt modalities, we aim to identify which modality leads to the best performance and evaluate the contribution of each modality to the overall performance.
Reasoning Experiment 2, the Text-only Setup (Instructional and Data Text)
In order to assess whether visual representations can be helpful, we first need to see the general problem-solving performance using pure textual prompt. In this experiment, we give MLLMs pure textual description of the problem containing the description of a specific question.
@@ 4.2 Ablation on Multimodal prompts
5
6In this graph:
7The distance between node 0 and 1 is 1,
8the distance between node 1 and 3 is 9,
9the distance between node 0 and 3 is 5,
10the distance between node 0 and 2 is 2,
11the distance between node 2 and 3 is 4.
@@
Reasoning Experiment 3, the Vision-rich-text Setup (Vision with Instructional and Data Text)
@@
Visual aids can be helpful in solving complex problems for humans. For instance, when dealing with a GCP, sketching a diagram of the graph and attempting different coloring schemes can aid in determining whether a valid solution exists. In this experiment, we aim to investigate whether providing MLLM with both textual and visual representations of a problem can improve their performance in solving it. To do so, we present the MLLMs with a complete textual description of the problem, followed by an accompanying image that depicts the problem visually. Our hypothesis is that the image can serve as a supplementary source of information to the textual description, potentially enhancing the model’s understanding of the problem and enabling it to generate more accurate solutions. By incorporating both textual and visual inputs, we aim to evaluate the MLLMs’ ability to integrate and process multimodal information in complementary with each other, which is a crucial aspect of their overall performance.
4.3 Evaluation Metrics
@@
Our evaluation methodology incorporates a suite of metrics to assess the reasoning ability of MLLMs. Building upon the metrics established by fan2023nphardeval , we expand our analysis to include a novel metric specifically tailored to quantify and rule out noise in analyzing models’ reasoning abilities. This sequence of metrics will start with the Recognition Accuracy (RA) metric, evaluating the recognition accuracy of the problems in the visual prompts, i.e., if the MLLMs can recognize what are in the questions. Among those questions that a MLLM can accurately recognize the input, we then use the Instruction-following Effective Rate (ER) metric, calculating the average chances that a MLLM cannot yield an output compatible with the rule-based answer parser. Finally, the Aggregated Accuracy (AA) will calculate the weighted accuracy of MLLMs’ answers correctness among those recognizable and parsable scenarios, aggregated with RA and ER. Figure 5 shows the sequence and relationship of these metrics.
Figure 5: Metrics Sequence
Recognition Accuracy (RA)
@@
Recognition Accuracy evaluates the MLLMs’ capability to accurately interpret the visual information presented in the prompts. This metric is crucial as it forms the foundation for any subsequent reasoning or decision-making processes the model undertakes. It is quantified by the ratio of prompts that are correctly recognized by the model to the total number of prompts presented, and is defined as:
@@
Figure 6: MLLM: a. Reasoning abilities performance excluding the effects of recognition and instruction following on figure+limited__\__text representation of questions. b. Instruction-following effective rate
Figure 7: LLM: a. Reasoning abilities performance excluding the effect of instruction following on pure textual description of questions. b. Instruction-following effective rate
Instruction-following Effective Rate (ER)
@@
Aggregated Accuracy (AA)
Aggregated Accuracy takes into account the correct recognition of prompts and the successful parsing of responses to evaluate the correctness of MLLMs’ answers. This metric is a weighted measure that integrates the Recognition Accuracy and the inverse of the Failure Rate:
@@
where $w_{i}$ represents the difficulty weight for the $i^{th}$ level, $A^{{}^{\prime}}_{i}$ represents the accuracy in recognizable and parsable questions at $i^{th}$ level, $RA_{i}$ is the Recognition Accuracy for the $i^{th}$ level, and $ER_{i}$ is the Instruction-following Effectiveness Rate for the $i^{th}$ level. The Aggregated Accuracy metric will be used as the main evaluation metric for the overall performance and the complexity-specific submetrics, demoted as $AA$ or $AA_{All}$, $AA_{NPH}$, $AA_{NPC}$, and $AA_{P}$.
@@
Figure 8: Reasoning abilities across models, complexity levels, and difficulty levels
5 Results
5.1 Reasoning Abilities of MLLMs
Close and Open Source Models
@@ 5.1 Reasoning Abilities of MLLMs
As Figure 1 and Figure 7 indicate, the comparison between close source and open source MLLMs is quite stark, with close source models exhibiting superior performance in all tasks, irrespective of complexity class. Specifically, demonstrated in Figure 8, the Gemini model consistently outperforms its counterparts, including the other close source model GPT-4V, across the board. A detailed inspection of the heatmap reveals that Gemini maintains a higher aggregated accuracy and a more effective Instruction-following rate than GPT-4V, which suggests that the proprietary nature in the Gemini model may contribute to enhanced reasoning performance. This is especially evident in the NP-hard category, where Gemini shows remarkable resilience in reasoning accuracy compared to GPT-4V, which exhibits a notable decline.
Complexity Classes
@@
As Figure 8 indicates, the reasoning capabilities of MLLMs are inversely proportional to the complexity of the tasks. On simpler P problems, these models show commendable performance. However, as the complexity escalates to NP-complete and further to NP-hard problems, a clear and expected downtrend in reasoning ability is observed. This trend is consistent across all MLLMs, signifying a universal challenge in tackling higher-order complexity with current multimodal reasoning architectures.
@@
In addition to the overall trend, there are two notable findings. First, none of the nine models are capable of solving two of the NP-hard problems, TSP and MSP. Second, the Otter model performs unexpectedly well on the TSP-D problem – all other models, including the close source ones are unable to solve any of the TSP-D problems. This may due to the specific training data relevant or similar to the TSP-D problem.
Task Difficulties
@@
When focusing on individual reasoning tasks and considering models that can at least address the simplest questions, we notice a degradation in performance in correlation with increasing question difficulty. As Figure 8 shows, models like Gemini display a high success rate on easier questions within tasks like GCP, but this success rate diminishes as the difficulty level of the questions increases. This pattern underscores the models’ limitations and suggests that even the most capable MLLMs struggle to maintain their reasoning prowess as task complexity intensifies.
MLLMs vs. LLMs
@@
Figure 6 and Figure 7 provide a direct comparison between the current top-performing models in MLLMs and LLMs, including both closed-source and open-source models. In LLMs, we selected the models GPT-4-turbo, GPT-3.5-turbo, Claude-2, Claude-instant, PaLM-2, MPT-30b, Mistral-7b, Phi-1.5, Phi-2, Qwen-14b, and Yi-34b fan2023nphardeval . The results indicate that MLLMs lag behind LLMs in reasoning tasks. Specifically, the aggregated accuracy in P problems is approximately 0.4 in LLMs and 0.2 in MLLMs. Furthermore, the decrease in aggregated accuracy is more pronounced in MLLMs compared to LLMs. While LLMs maintain a relatively consistent weighted accuracy in NP-complete and P problems, MLLMs’ weighted accuracy decreases dramatically by half on average, from approximately 0.2 to 0.1. These findings highlight the need for further research and development to improve the reasoning abilities of MLLMs.
@@
The results of our study indicate that the integration of multimodal data processing in MLLMs does not necessarily lead to improved reasoning capabilities. In fact, our findings reveal that the current development of MLLMs falls short in comparison to LLMs, with significantly weaker performance in reasoning tasks. This highlights the need for further research and development efforts aimed at enhancing the reasoning abilities of MLLMs.
5.2 Impact of Vision and Text Input
Figure 9: Reasoning abilities across prompt types

### v3

@@ 3.2.1. Graph Data Transformation
Figures in algorithmic questions such as SPP, GCP_\_D, TSP, and TSP_\_D are transformed from text similarly: these figures include graphs containing nodes and edges; in the case of TSP and TSP_\_D, the weights on the edges are also included.
3.2.2. Linear Data Transformation
Similarly, linear data problems like the Knapsack Problem (see Figure 2) are presented with both a prompt and a visual representation. In order to strictly align the input content across modalities, our generation scripts ensure that the visual representation is a neutral and information-preserving rendering of the textual data. For instance, in the Knapsack problem, we use Python to create blocks with sizes and labels that precisely correspond to the weights and IDs of the items described in the text. This programmatic approach is designed to ensure informational equivalence and avoid one modality inadvertently containing more prompt information than the other.
@@ 3.2.2. Linear Data Transformation
For algorithmic problems such as EDP and BSP, we do not employ a specialized visualization method. Instead, we simply display the relevant string of characters or string of numbers in the image.
In summary, we employ various visualization functions to transform textual descriptions into visual representations for different algorithmic problems. In the prompt, we provide a combination of textual and visual information to the model, including an instructional prompt and a general introduction of the algorithm, along with an image that represents the specific problem to be solved. This approach is demonstrated in the previous examples for GCP and KSP. By providing both textual and visual information, we aim to evaluate the reasoning abilities of LVLMs in handling complex problems that require both visual and textual understanding.
3.3. Dynamic Benchmark Mechanism
A core feature of NPHardEval4V is its dynamism, designed to prevent benchmark overfitting and provide a continuous challenge for evolving models. This dynamism is implemented through a programmatic generation pipeline.
Implementation of Dynamism: The benchmark is powered by a suite of algorithms that can generate new, unique problem instances on demand for each task (e.g., KSP, TSP). For each new version of the benchmark, our generator program is executed to create a fresh dataset of 900 questions (9 tasks × 10 difficulty levels × 10 questions per level). This ensures that models cannot simply memorize solutions from previous versions. The generation process guarantees that while the underlying logic of the task remains the same, the specific parameters (e.g., graph structures, item weights, city coordinates) are novel, ensuring the unpredictability of samples.
Version Control and Update Frequency: To maintain consistency and track progress over time, each generated dataset is versioned. We recommend a bi-annual or annual update frequency, which strikes a balance between keeping pace with rapid model development and providing a stable target for the research community.
4. Experimental Setting
To address the outlined research questions, our experimental setting systematically assesses the performance of various LVLMs in recognition and reasoning tasks and investigates the impact of vision and text inputs on LVLMs’ performance. The models we test on are presented in Table 2. We detail the experimental setup corresponding to each research question, describing the models evaluated, the nature of the tasks, the methodology of the performance assessment, and the structure of the benchmark.
@@ 4. Experimental Setting
GPT-4o (OpenAI, 2024) Unknown Close source
Gemini-2.0-flash-exp (Google, 2023) Unknown Close source
claude-3.7-sonnet (Anthropic, 2025) Unknown Close source
Llama-3.2-vision (Huang et al., 2024) 90 Open source
InternVL3 (Zhu et al., 2025) 78 Open source
Llava-qwen2 (Li et al., 2024) 72 Open source
Qwen2.5-VL (Team, 2025) 72 Open source
Qwen2-VL (Wang et al., 2024a) 72 Open source
Nvlm-D (Dai et al., 2024) 72 Open source
Ovis2 (Lu et al., 2024) 34 Open source
DeepSeekVL2 (Wu et al., 2024) 27 Open source
Gemma (Team et al., 2025) 27 Open source
GLM-4V (GLM et al., 2024) 9 Open source
Ola (Liu et al., 2025) 7 Open source
Point1.5-qwen2.5 (Liu et al., 2024) 7 Open source
Table 2. LVLMs metadata
4.1. Recognition and Reasoning Performance Evaluation
@@
In the initial experiment, the objective is to assess the performance of Large Vision-Language Models (LVLMs) in terms of recognition and reasoning across a range of benchmark problems. To this end, zero-shot prompts are utilized to evaluate the inherent recognition and reasoning capabilities of each model. The performance of the models is then analyzed and compared across the benchmark problems to draw conclusions about their relative strengths and weaknesses in recognition and reasoning tasks.
4.1.1. Recognition Experiment
@@
In order to precisely assess the capacity for reasoning in Large Vision-Language Models (LVLMs), it is essential to first comprehend their proficiency in image recognition, as accurate reasoning can only be established upon correct recognition. Therefore, an evaluation of recognition capabilities is initially conducted. For each question, both a visual representation of the description and a textual representation are provided. Subsequently, the MLLM is presented with both the visual and textual representations and is asked to determine whether they correspond to the same question. To mitigate the influence of randomness, the model is prompted five times, each time with a distinct random seed. The mean and variance of the model’s responses are then calculated, and it is concluded that the model demonstrates recognition capabilities if it correctly identifies the correspondence between the visual and textual representations more than half of the time, i.e., answering ’yes’ three or more times. As an preprocess for analyzing reasoning ability, recognition experiment reuslts and analysis are provided in Appendix C
@@
4.1.2. Reasoning Experiment 1, the Default Setup (Vision with Instructional Text)
@@
This experiment evaluates the overall performance of MLLM on a benchmark consisting of various tasks. In the setting, a textual prompt is provided that includes a general introduction to the question and the format in which the answer should be given. Subsequently, an image is presented that pertains to the specific question being asked. The models are then tasked with processing both the textual and visual information in order to generate accurate responses. The performance of the LVLMs is evaluated based on their ability to correctly answer the questions, taking into account both their recognition and reasoning capabilities. The results of this experiment provide insights into the overall effectiveness of the models in handling multimodal tasks and their potential applications in real-world scenarios.
@@
Figure 3. Image of an Short Path Problem example
In order to accurately evaluate the reasoning performance, it is necessary to isolate the specific aspect of reasoning ability from other factors that may influence the overall performance: we employ a filtering process that removes datapoints where the model fails to recognize the given input or fails to provide a parsable result, which indicates a failure in instruction-following. After filtering out these two factors, we can obtain a more accurate picture of the LVLMs’ pure reasoning ability.
4.2. Ablation on Multimodal prompts
The goal of this experiment is to investigate the impact of prompt modality on the reasoning performance of LVLMs. Specifically, we aim to assess whether the inclusion of visual or textual inputs can enhance the models’ problem-solving abilities. To achieve this, we conduct an ablation study, varying the modality of the prompts provided to the LVLMs by pure textual prompt and textual + visual prompt. By comparing the models’ performance across different prompt modalities, we aim to identify which modality leads to the best performance and evaluate the contribution of each modality to the overall performance.
4.2.1. Reasoning Experiment 2, the Text-only Setup (Instructional and Data Text)
In order to assess whether visual representations can be helpful, we first need to see the general problem-solving performance using pure textual prompt. In this experiment, we give LVLMs pure textual description of the problem containing the description of a specific question.
@@ 4.2.1. Reasoning Experiment 2, the Text-only Setup (Instructional and Data Text)
4Your output should be enclosed within
5<root></root> tags. Include your reasoning
6in <reasoning></reasoning> tags and the final
7path and total distance in <final_answer>
8</final_answer> tags, like <final_answer>
9{’Path’: START->...->END’, TotalDistance’:
10INT_TOTAL_DISTANCE’}</final_answer>.
11
12In this graph:
13The distance between node 0 and 1 is 1,
14the distance between node 1 and 3 is 9,
15the distance between node 0 and 3 is 5,
16the distance between node 0 and 2 is 2,
17the distance between node 2 and 3 is 4.
@@
4.2.2. Reasoning Experiment 3, the Vision-rich-text Setup (Vision with Instructional and Data Text)
@@
Visual aids can be helpful in solving complex problems for humans. For instance, when dealing with a GCP, sketching a diagram of the graph and attempting different coloring schemes can aid in determining whether a valid solution exists. In this experiment, we aim to investigate whether providing MLLM with both textual and visual representations of a problem can improve their performance in solving it. To do so, we present the LVLMs with a complete textual description of the problem, followed by an accompanying image that depicts the problem visually. Our hypothesis is that the image can serve as a supplementary source of information to the textual description, potentially enhancing the model’s understanding of the problem and enabling it to generate more accurate solutions. By incorporating both textual and visual inputs, we aim to evaluate the LVLMs’ ability to integrate and process multimodal information in complementary with each other, which is a crucial aspect of their overall performance.
4.3. Evaluation Metrics
@@
Our evaluation methodology incorporates a suite of metrics to assess the reasoning ability of LVLMs. Building upon the metrics established by (Fan et al., 2024a), we expand our analysis to include a novel metric specifically tailored to quantify and rule out noise in analyzing models’ reasoning abilities. This sequence of metrics will start with the Recognition Accuracy (RA) metric, evaluating the recognition accuracy of the problems in the visual prompts, i.e., if the LVLMs can recognize what are in the questions. Among those questions that a MLLM can accurately recognize the input, we then use the Instruction-following Effective Rate (ER) metric, calculating the average chances that a MLLM cannot yield an output compatible with the rule-based answer parser. Finally, the Aggregated Accuracy (AA) will calculate the weighted accuracy of LVLMs’ answers correctness among those recognizable and parsable scenarios, aggregated with RA and ER. Figure 4 and Figure 5 shows the sequence and relationship of these metrics.
Figure 4. Multimodal Large Language Models’s performance on recognition (RA), Instruction-following (ER), and reasoning (AA) on polynomial time (P), NP-complete, and NP-hard problems. All scores range from 0 to 1, with higher values indicating better performance. The a plot shows closed-source models (Claude-3.7, Gemini-2, GPT-4o), while the b and c plots show two groups of open-source models. Larger radar areas indicate stronger overall capabilities, while smaller or skewed shapes reveal specific weaknesses, particularly in handling harder reasoning tasks.
Figure 5. Metrics Sequence
4.3.1. Recognition Accuracy (RA)
@@
Recognition Accuracy evaluates the LVLMs’ capability to accurately interpret the visual information presented in the prompts. This metric is crucial as it forms the foundation for any subsequent reasoning or decision-making processes the model undertakes. It is quantified by the ratio of prompts that are correctly recognized by the model to the total number of prompts presented, and is defined as:
@@
Figure 6. MLLM: a. Reasoning abilities performance excluding the effects of recognition and instruction following on figure+limited_\_text representation of questions. b. Instruction-following effective rate. LLM: c. Reasoning abilities performance excluding the effect of instruction following on pure textual description of questions. d. Instruction-following effective rate
4.3.2. Instruction-following Effective Rate (ER)
@@
4.3.3. Aggregated Accuracy (AA)
Aggregated Accuracy takes into account the correct recognition of prompts and the successful parsing of responses to evaluate the correctness of LVLMs’ answers. This metric is a weighted measure that integrates the Recognition Accuracy and the inverse of the Failure Rate:
@@ 4.3.3. Aggregated Accuracy (AA)
where $w_{i}$ represents the difficulty weight for the $i^{th}$ level, $A^{{}^{\prime}}_{i}$ represents the accuracy in recognizable and parsable questions at $i^{th}$ level, $RA_{i}$ is the Recognition Accuracy for the $i^{th}$ level, and $ER_{i}$ is the Instruction-following Effectiveness Rate for the $i^{th}$ level. The Aggregated Accuracy metric will be used as the main evaluation metric for the overall performance and the complexity-specific submetrics, demoted as $AA$ or $A_{All}, AA$$_{NPH}, AA$$A_{NPC}$, and $AA_{P}$.
@@
Figure 7. Reasoning abilities across models, complexity levels, and difficulty levels
5. Results
This section presents the findings from our experiments using the NPHardEval4V benchmark. We evaluate and compare the reasoning abilities of various Large Vision-Language Models (LVLMs), considering factors such as model type (open vs. closed source), task complexity, question difficulty, and the comparison between LVLMs and traditional LLMs.
5.1. Reasoning Abilities of LVLMs
@@
To better understand the reasoning potential of LVLMs, we analyze their performance across several axes, starting with the difference between open-source and closed-source models. As Figure 4 and Figure 6 indicate, there is a huge difference in performance between close source and open source LVLMs, with close source models exhibiting superior performance in all tasks, irrespective of complexity class. Specifically, demonstrated in Figure 7, the Claude-3.7 model consistently outperforms its counterparts, including the other close source model Gemini and GPT-4o, across the board. A detailed inspection of the heatmap reveals that Claude-3.7 maintains a higher aggregated accuracy and a more effective Instruction-following rate than Gemini, which suggests that the proprietary nature in the Claude-3.7 model may contribute to enhanced reasoning performance. This is especially evident in the NP-hard category, where Claude-3.7 shows remarkable resilience in reasoning accuracy compared to Gemini, which exhibits a notable decline.
@@
As Figure 7 indicates, the reasoning capabilities of LVLMs are inversely proportional to the complexity of the tasks. On simpler P problems, these models show commendable performance. However, as the complexity escalates to NP-complete and further to NP-hard problems, a clear and expected downtrend in reasoning ability is observed. This trend is consistent across all LVLMs, signifying a universal challenge in tackling higher-order complexity with current multimodal reasoning architectures.
@@
In addition to the overall trend, there are two notable findings. First, none of the fifteen models are capable of solving two of the NP-hard problems, TSP and MSP. Second, the Qwen2.5-vl, Ovis and Internvl model performs unexpectedly well on the TSP-D problem – close to or even better than closed-source models. This may due to the specific training data relevant or similar to the TSP-D problem.
@@
When focusing on individual reasoning tasks and considering models that can at least address the simplest questions, we notice a degradation in performance in correlation with increasing question difficulty. As Figure 7 shows, models like Gemini display a high success rate on easier questions within tasks like GCP, but this success rate diminishes as the difficulty level of the questions increases. This pattern underscores the models’ limitations and suggests that even the most capable LVLMs struggle to maintain their reasoning prowess as task complexity intensifies.
@@
Figure 6 provide a direct comparison between the current top-performing models in LVLMs and LLMs, including both closed-source and open-source models. For LLMs, we selected the models GPT-4-turbo, GPT-3.5-turbo, Claude-2, Claude-instant, PaLM-2, MPT-30b, Mistral-7b, Phi-1.5, Phi-2, Qwen-14b, and Yi-34b (Fan et al., 2024a). The results indicate that LLMs lag behind LVLMs in reasoning tasks. Specifically, the aggregated accuracy in P problems is approximately 0.4 in LLMs and 0.48 in LVLMs. However, the decrease in aggregated accuracy is more pronounced in LVLMs compared to LLMs. While LLMs maintain a relatively consistent weighted accuracy in NP-complete and P problems, LVLMs’ weighted accuracy decreases 35.42%, from approximately 0.48 to 0.31. These findings highlight the need for further research and development to improve the reasoning abilities of LVLMs.
@@
The results of our study indicate that the integration of multimodal data processing in LVLMs does not necessarily lead to improved reasoning capabilities. In fact, our findings reveal that the current development of LVLMs falls short in comparison to LLMs, with significantly weaker performance in reasoning tasks. This highlights the need for further research and development efforts aimed at enhancing the reasoning abilities of LVLMs.
5.2. Impact of Vision and Text Input
Figure 8. Reasoning abilities across prompt types